A train travels 250 km at the speed of 60 km/hr and arrives ten minutes late later than schedule. At what speed (in km/hr) should it travel to arrive exactly on time ?
Answer by mananth(16946)

let the original speed be x
t=d/r
time taken - scheduled time = 1/4 hour

(250/60)-(250/x)=1/4
LCD = 60 x
250x-250*60=60x/4
multiply by 4
1000x-1000*60=60x
960x=60000
/960
x= 62.5 km/h